Circular No. 5737

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

SUPERNOVA 1993J IN NGC 3031 M. Richmond, Princeton University, provides the following J2000.0 po-

sition of the supernova, reduced from the Treffers CCD images (IAUC 5731) using the Guide Star Catalogue: α = 9^{h}55^{m}25^{s}.00, δ = +69^{o}01′13′′.3. The corresponding end figures for the position by Hartwick et al. (IAUC 5731) are 25^{s}.00, 13′′.7. Richmond gives the offset from the mag 14 foreground star mentioned on IAUC 5731 (= GSC 4383.0340) as 7′′.4 east, 24′′.3 north (7′′.2 east, 24′′.3 north at 1950.0).

W. van Driel, Institute of Astronomy, University of Tokyo, reports: “On Mar. 25.60 UT an R-band CCD image of M81 was obtained with the 1.05-m Schmidt telescope at the Kiso Observatory by K.-I. Wakamatsu and myself. The 5-min exposure was made through a thin cloud layer. We can not see the supernova in this image, though we can see the mag 17 star 50′′ due north of GSC 4383.0340. On Mar. 31 we took R and V band images of the supernova for comparison. The Kiso CCD is 1018 × 1000 pixels of 0′′.75 square each, giving a field-of-view of ∼ 12′.7 × 12′.5. We have performed the bias substraction and flatfielding of the images.”

A. V. Filippenko, University of California at Berkeley, notes: “The V magnitude and colors (B − V, V − R) of the starlike object reported on IAUC 5736 at the position of SN 1993J are consistent with those of an unreddened or slightly reddened K0 Ia supergiant at the distance of NGC 3031 (distance modulus 27.5 mag). On page 16 of the Atlas of Galaxies by Sandage and Bedke (1988) the object is clearly visible as a stellar or semistellar condensation in the middle of a spiral arm. An extended progenitor (red supergiant, rather than blue supergiant like that of SN 1987A) is consistent with the fact that SN 1993J is not subluminous in B.”

J. C. Wheeler, University of Texas at Austin, reports: “E. Barker obtained an essentially featureless spectrum from 300 to 840 nm with the es2 spectrograph on the 2.1-m telescope at the McDonald Observatory on Mar. 31.2 UT. D. Lester, N. Gaffney and B. Smith obtained spectra with the McDonald Observatory infrared spectrometer on the 2.7-m telescope, covering the J and K bands at low (R = 100-200) resolution. These spectra are also basically featureless with no evidence for Brackett, Paschen or He lines. Approximate photometry is K = 10.6, J = 10.9 on Mar. 31.2.”

Visual magnitude estimate by M. V. Zanotta and S. Pesci (Milan, Italy): Mar. 31.84 UT, 10.5.

Corrigendum. IAUC 5733: for FWZI ∼ 70 nm read FWZI ∼ 37 nm.

1993 April 1 Brian G. Marsden